I have to write an equation in standard form of the line that passes through the two points:
(4,-7),(5,-1)
thanks :)

Answer by Alan3354(69443)

write an equation in standard form of the line that passes through the two points:
(4,-7),(5,-1)
=====================================
Do it like this one:
-------
write an equation in standard form of the line that passes through the two points:
(-3, -3),(7,2)
-------------
Find the slope, m
m = diffy/diffx = (-3-2)/(-3-7) = -5/-10
m = 1/2
----------
Use y = mx + b and either point to find b
2 = (1/2)*7 + b
b = -3/2
---------
eqn is y = mx + b
y = (1/2)x - 3/2 (slope-intercept form)
------
x - 2y = 3 (standard form)